OpenStudy (anonymous):

You pay 3.3 cents per minute for long-distance phone calls on your phone. Your phone bill is $3.40. Write and solve an equation to find the number of minutes m of long-distance calls you made. A. ; 103 minutes B. ; 11 minutes C. ; 83 minutes D. ; 10 minutes

OpenStudy (anonymous):

3.3*x=340 x=340/3.3 x=103.030303.... It's A